OpenStudy (anonymous):

PLEASE HELP whats the equation for slope 3/4, passes through (4,-7)

OpenStudy (anonymous):

y-y1=m(x-x1) y--7=3/4(x-4) y+7=3/4x-3 y=3/4x+4

OpenStudy (anonymous):

wait never mind its y=3/4x-10

OpenStudy (anonymous):

thank you! @S.Kumar

OpenStudy (anonymous):

welcome:)